# Particle production and hadronization temperature of the massive Schwinger model

### v1

@@ I Introduction
This suggests that there is another mechanism at play and we have previously discussed one based on entanglement between spatial regions in a QCD string [6, 7, 8]. Here we want to continue this line of thought and investigate a more complete model for the quantum dynamics of hadronization.
Computing the hadronization process in the collision from first principles is challenging for several reasons. First, the process is non-perturbative [9], meaning that it involves QCD dynamics at large distances where the coupling is strong. Moreover, describing the process accurately requires taking into account the dynamics in time of the quark-antiquark expanding string as it breaks into hadrons during particle production.
Several theoretical models have been proposed [10], but apparently none can account for all the observed features. The Lund model [11, 12] is based on expanding QCD strings from which tunneling processes produce hadrons and resonances via the Schwinger mechanism [13]. The standard implementation of the Lund model is the PHYTHIA event generator [14, 15, 16, 17] for which it seems difficult to explain the thermal-like features seen in experimental data without additional modifications [17].
As explained above, the full problem of particle production in QCD is challenging. Therefore, one can resort to effective models that are simpler to study and can still explain several aspects of the original model. The Schwinger model is quantum electrodynamics in 1+1 space-time dimensions (QED1+1[18, 19, 20, 21]. Despite being a $U(1)$ abelian gauge theory, it shares several essential features with QCD, which is a $SU(3)$ gauge theory, making it an effective toy model for simulating some QCD characteristics [22]. QED1+1 includes the Higgs mechanism, charge screening, “quark” confinement [23, 24], string breaking [11, 25, 26, 27, 28, 29, 30, 31], and spontaneous chiral symmetry breaking, as well as topological vacua. Furthermore, transverse and longitudinal degrees of freedom naturally separate at high energies, thus justifying the use of a dimensionally reduced theory.
@@
Recently, there has been much interest in low-dimensional dynamics with the ultimate goal of simulating gauge theories using quantum simulators [36]. Various works have proposed the analog simulation of the massive Schwinger model using ultracold quantum gases [37, 38, 39, 40, 41], and the quantum simulation of its non-perturbative aspects using tensor networks [42], and spin chains [43]. We explore a possible experimental realization of this model using ultracold atoms, in particular, two tunnel-coupled Bose-Einstein Condensates (BEC) with a modulated linear interaction.
@@ II.2 Bosonized action
The theory in Eq. (1) is alternatively described by the microscopic action
with the scalar potential
$u_{\Lambda}=\frac{\exp(\gamma)}{2\pi}\Lambda m,$ (6)
II.3 Renormalization
The bosonized theory in Eq. (4) is an interacting quantum field theory subject to renormalization. A detailed discussion of renormalization in the functional formulation, addressing states with zero and non-zero temperature, can be found in Ref. [22]. The path integral systematically includes quantum fluctuations into correlation functions when moving from UV to IR, where the scale-dependent fluctuations enter corresponding coupling constants of the scale-dependent effective action $\Gamma_{k}$. In the limit $k\gg\Lambda$ all quantum fluctuations are suppressed, such that $\Gamma_{k}[\phi]=S_{\Lambda}[\phi]$. On the other side, for $k\to 0$, the flowing action would approach the full quantum effective action $\Gamma_{k=0}[\phi]=\Gamma[\phi]$, see Ref. [22] for a more detailed discussion. The regulator function is chosen such that $\Gamma_{k}$ smoothly interpolates between the microscopic action $S_{\Lambda}$ for $k=\Lambda$ and the macroscopic action $\Gamma$ for $k=0$. We assume that $\Gamma_{k}[\phi]$ is of the same form as Eq. (4) with potential term as in Eq. (5), except for the replacement $u_{\Lambda}\to u_{k}$, whereas the derivative term keeps its bare form. For the purpose of the present work, we shall be satisfied with this simplified approach where only a single parameter, the coupling $u_{\Lambda}$, gets replaced by the renormalized coupling $u$. There is, in fact, a good reason why $M$ does not need to be renormalized, which is explained below. The $k$-dependent effective potential satisfies the exact equation [45]
$\begin{split}\partial_{k}V_{k}(\phi)=&\frac{1}{2}\int\limits_{p^{2}<\Lambda^{2}}\frac{\mathrm{d}^{2}p}{(2\pi)^{2}}\frac{2k}{p^{2}+k^{2}+V_{k}^{(2)}(\phi)}\\ =&\frac{k}{4\pi}\ln\left(\frac{\Lambda^{2}+k^{2}+V_{k}^{(2)}(\phi)}{k^{2}+V_{k}^{(2)}(\phi)}\right),\end{split}$ (7)
with
To obtain a flow equation for $u_{k}$, we use the inverse Fourier expansion scheme
This can be integrated with respect to $k$ to yield
$u=u_{0}=u_{\Lambda}\frac{M}{\Lambda}.$ (12)
In summary, the renormalized action reads in our approximation
@@ II.3 Renormalization
The potential in Eq. (14) was also found in Ref. [23]. We note that it is a renormalized potential in the sense that some quantum fluctuations have been taken into account through the renormalization of $u$. However, it cannot be seen as the effective potential corresponding to full quantum effective action, which would have to be convex as a Legendre transform. In the following, we also work with the dimensionless ratio
@@
II.4 Potential
Figure 1: Potential of the massive Schwinger model $V(\phi)$ (Eq. (14)) for $\theta=0$ (left panel) and $\theta=\pi$ (right panel). The different colors represent different dimensionless coupling strengths $\kappa$ (Eq. (16)). The dots indicate the initial condition for the field according to Eq. (20).
@@ II.4 Potential
II.5 Weak coupling limit
In the weak coupling limit $m\gg q$, the quadratic term $\sim M^{2}$ in the potential in Eq. (14) (or, equivalently for the microscopic potential in Eq. (5)) is subleading compared to the cosine term and one has effectively a kind of sine-Gordon theory [49]. It is instructive to discuss this limit where intuition from perturbative QED can be applied.
@@ III Dynamics of expanding strings
Let us now turn to the problem of an expanding QCD string in the reduced model, where the dynamics happens in $d=1+1$ dimensions and massive QED is employed instead of QCD. We start from a quark-antiquark pair with very high energy, where the quark and anti-quark move in opposite directions with the speed of light. The electric field that forms between them will be modeled as a coherent scalar field in the bosonized description and we will study its evolution, as well as the excitations around it.
@@ III.1 Coordinate system
Expressed in Bjorken coordinates, the equation of motion for the scalar field as obtained from the variation of Eq. (13) reads (using the ratio in Eq. (16))
@@ III.2 Background field evolution
In order to fully understand the evolution of the coherent field $\Phi(\tau)$ also at late times, we numerically solve the non-linear equation of motion, Eq. (19), using the fourth-order Runge-Kutta method. Fig. 2 illustrates the behavior of $\Phi(\tau)$ for various choices of the ratio $\kappa$ and the vacuum angles $\theta=0$ as well as $\theta=\pi$.
@@
Note that depending on the value of $\kappa$ there are qualitative differences in the dynamics of the background field. For very small values $\kappa\ll 1$, corresponding essentially to small fermion mass $m\ll q$, the potential has only a single global minimum which is then approached by the classical evolution equation in the asymptotic long time limit. In contrast, when $\kappa$ is larger, the background field can be trapped in a local minimum, which corresponds to a metastable state.
In section V, we will discuss the production of quantum excitations around the coherent background field solution $\Phi(\tau)$. However, before that we first address the metastable background solutions and the decay of such false vacuum states through field tunneling.
@@ IV False vacuum decay
It is an interesting problem to follow the time evolution in real time, within the future light cone of the breaking point at the coordinate origin. This leads to a very similar problem to the one discussed in the previous section. Directly on the light cone, initial conditions are set through the bounce solution, but everything in the future of that is to be computed dynamically.
@@
Displayed on the lower plots of Fig. 3 we show the resulting background field $\Phi$ evolution for the new initial conditions (for $\theta=0$ in the left panel and $\theta=\pi$ in the right panel). Again, the field shows damped oscillations, now around the true vacuum.
@@
Combining both solutions, the analytically continued bounce solution for $|x|>t$ and the solution to classical evolution equations for $t>|x|$ is shown in Fig. 4. The field decays from the false vacuum outside the light cone to the true one inside it, showing damped oscillations very similar to those obtained in section III.2.
@@ V Particle production
Now that we have obtained non-trivial evolving background field configurations in two closely related scenarios, it is interesting to study quantum fluctuations around them. Field excitations can be seen as particles, and in this section we aim to compute the resulting particle production.
@@ V.1 Adding quantum fluctuations
such that the equal-time canonical commutation relation reads with the conjugate momentum field $\Pi(\tau,\eta)=\delta\Gamma/\delta(\partial_{\tau}\phi(\tau,\eta))=\tau\partial_{\tau}\phi(\tau,\eta)$:
@@ V.2 Initial choice of mode functions
Another interesting state to be considered is a coherent state that basically agrees with the vacuum state, but where the expectation value is shifted, $\Phi_{\text{vac}}\to\Phi(\tau_{0})$. We denote this state by $|\Omega,\Phi(\tau_{0})\rangle$ with $\Phi(\tau_{0})$ the field expectation value $\Phi(\tau)$ at the initial time $\tau=\tau_{0}$. For the scenario discussed in section III this initial value is given in Eq. (20), while for the scenario discussed in section IV it is specified in Eq. (26). For the mode functions in this state, we use at very early times $\tau\to 0^{+}$ the form given in Eq. (32) without any modification. Of course, deviations will occur at later times $\tau>0$, see below.
The “shifted vacuum” coherent state specified through this prescription is only one out of many possible quantum states one could consider at $\tau=0^{+}$. In particular for the bounce scenario discussed in section IV one may expect that one could do better and actually determine the quantum state of fluctuations after the bounce possibly through analytic continuation from a calculation in Euclidean space, but we leave this for future work.
Starting with the initial conditions (32) for very early times, we use the differential equations (29) to find solutions for later times. Figure 5 shows the real and imaginary part of the mode functions obtained in this way for different choices of the coupling constant $\kappa$ and the rapidity wavenumber $k$. In the right panel of Fig. 6, we show the growth in time of the occupation numbers (see Eq. (38) below for their definition) for several wavenumber $k$ for $\theta=0,\kappa=4.0$. The occupation numbers exhibit exponential growth for a time window, after which their value saturates (see Ref. [60] for a detailed discussion).
@@ V.3 Asymptotic particle number
As a consequence of the overall expansion and dilution of energy, the background field $\Phi(\tau)$ eventually approaches a constant asymptotic value for very late Bjorken times, $\Phi(\tau)\to\Phi_{\text{asym}}$ for $\tau\to\infty$. We concentrate on situations where this value agrees with the global minimum, $\Phi_{\text{asym}}=\Phi_{\text{vac}}$. (For $\theta=\pi$, there are two degenerate global minima, and $\Phi(\tau)$ dynamically selects one.)
@@
Using the usual Bogoliubov theory arguments (see, for example, Ref. [44, 61]), one can obtain the following expression for the occupation number of the asymptotic state
@@
Because $f_{k}^{\text{asym}}(\tau)$ is only a proper solution to the mode equation (29) at asymptotically late time, the expression in Eq. (38) only has strict physical meaning in the limit $M\tau\gg 1$. The left panel of Fig. 6 shows the particle number $n_{k}$ against the wave vector $k$ at different times $\tau$ for the choice of parameters $\theta=0$ and $\kappa=4.0$ (left panel). One observes that around $M\tau=510$ the convergence with $M\tau$ is approaching, at least for some choices of $\theta$, $\kappa$, and $k$. The spectrum displays a peak at $k\approx 0$. This is the expected behavior: excitations with small energy can be produced, while for large $k$, the kinetic term $\sim k^{2}$ in Eq. (29) dominates over the time-dependent term in the second line and particle production is suppressed.
@@ V.4 Total particle number and temperature
which are of the same order of magnitude as the hadronization temperature $T_{H}\approx 160$--$170$ MeV [4]. The parameter $\theta=\pi$ seems to be closer to the actual result. However, $\theta=0$ is of the same order of magnitude, indicating that other angles may yield comparable results. Overall, it can be assumed that the numerical outcome is relatively consistent across various angles and different choices of the initialization of the fluctuations and that this result remains robust. Some discrepancies in the results are expected since QED1+1 is only a toy model for QCD.
The above results have been obtained for fluctuations around a decaying false vacuum as described in section IV. The calculation can also be done for fluctuations around the background solution described in section III with a similar result, but temperatures given by $T_{\theta=0}\approx 1\text{ GeV},T_{\theta=\pi}\approx 2\text{ GeV}.$
@@ VI.1 Experimental setup
The two components in the system are represented by the field operators $\psi_{1}$ and $\psi_{2}$. Both have the same mass $m$, and their interactions are determined by two coupling constants, namely $h$, the coupling of each component with itself, and $h_{c}$, which is related to an inter-component interaction. The external potential of the trap in which the atoms are confined has been omitted here. Additionally, an external field, e.g., a radio frequency field, induces the transition between the two states with a transition rate $\nu$. Furthermore, the transition rate $\nu_{0}$ is modulated with radio frequency $\omega$, the amplitude of the modulation being $\nu(t)=\nu_{0}+\delta\omega\cos(\omega t)$. To bring the Hamiltonian (44) into a more convenient form, the fields are expressed in terms of their densities $\rho_{j}=\psi_{j}^{\dagger}\psi_{j}$ and phases $\vartheta_{j}$,
@@
Additionally, we can switch to the set of variables that consists of the mean $\rho$ and the relative density $\epsilon$ defined as $\rho=(\rho_{1}+\rho_{2})/2,\,\epsilon=(\rho_{2}-\rho_{1})/2,$ and the sum $\zeta$ and the difference $\phi$ of the phases $\zeta=\vartheta_{1}+\vartheta_{2},\,\phi=\vartheta_{2}-\vartheta_{1}.$ The relative phase can be measured with an interferometer [67]. We will focus on the relative phase $\phi$ in the following. This should not be confused with the abovementioned scalar field $\phi$. Moreover, the frequency variation $\omega$ is assumed to be large in comparison to the timescales of the system, specifically $\omega_{0}=2\sqrt{\nu_{0}h\bar{n}}$ where $\bar{n}$ is the expectation value of the mean particle density. This allows time averaging and gives the final effective Lagrangian after integrating out the densities $\rho$ and $\epsilon$

### v2

# Particle production and hadronization temperature in the massive Schwinger model

@@ I Introduction
This suggests that there is another mechanism at play, and we have previously discussed one based on entanglement between spatial regions in a QCD string [6, 7, 8]. Here, we want to continue this line of thought and investigate a more complete model for the quantum dynamics of hadronization.
Computing the hadronization process in the collision from first principles is challenging for several reasons. First, the process is nonperturbative [9], meaning that it involves QCD dynamics at large distances where the coupling is strong. Moreover, describing the process accurately requires taking into account the dynamics in time of the quark-antiquark expanding string as it breaks into hadrons during particle production.
Several theoretical models have been proposed [10], but apparently, none can account for all the observed features. The Lund model [11, 12] is based on expanding QCD strings from which tunneling processes produce hadrons and resonances via the Schwinger mechanism [13]. The standard implementation of the Lund model is the PYTHIA event generator [14, 15, 16, 17] for which it seems difficult to explain the thermal-like features seen in experimental data without additional modifications [17].
As explained above, the full problem of particle production in QCD is challenging. Therefore, one can resort to effective models that are simpler to study and can still explain several aspects of the original model. The Schwinger model is quantum electrodynamics in 1+1 space-time dimensions (QED1+1[18, 19, 20, 21]. Despite being a $U(1)$ Abelian gauge theory, it shares several essential features with QCD, which is a $SU(3)$ gauge theory, making it an effective toy model for simulating some QCD characteristics [22, 23]. QED1+1 includes the Higgs mechanism, charge screening, “quark” confinement [24, 25], string breaking [11, 26, 27, 28, 29, 30, 31, 32], and spontaneous chiral symmetry breaking, as well as topological vacua. Furthermore, transverse and longitudinal degrees of freedom naturally separate at high energies, thus justifying the use of a dimensionally reduced theory.
@@
Recently, there has been much interest in low-dimensional dynamics with the ultimate goal of simulating gauge theories using quantum simulators [37, 38, 39, 40]. Various works have proposed the analog simulation of the massive Schwinger model using ultracold quantum gases [41, 42, 43, 44, 45], and the quantum simulation of its nonperturbative aspects using tensor networks [46], and spin chains [47]. We explore a possible experimental realization of this model using ultracold atoms, in particular, two tunnel-coupled Bose-Einstein condensates (BEC) with a modulated linear interaction.
@@ II.2 Bosonized action
We discuss the bosonized formulation and its renormalization in Appendix B. It is shown there that a renormalized effective action can be written in the form
@@
The potential in Eq. (5) was also found in Ref. [24]. We note that it is a renormalized potential in the sense that some quantum fluctuations have been taken into account through the renormalization of $u$. However, it cannot be seen as the effective potential corresponding to full quantum effective action, which would have to be convex as a Legendre transform.
In the following, we also work with the dimensionless ratio
@@
II.3 Potential
Figure 1: Potential of the massive Schwinger model $V(\phi)$ [Eq. (5)] for $\theta=0$ (left panel) and $\theta=\pi$ (right panel). The different colors represent different dimensionless coupling strengths $\kappa$ [Eq. (7)]. The dots indicate the initial condition for the field according to Eq. (11).
@@ II.3 Potential
II.4 Weak coupling limit
In the weak coupling limit $m\gg q$, the quadratic term $\sim$ $M^{2}$ in the potential in Eq. (5) [or, equivalently for the microscopic potential in Eq. (45)] is subleading compared to the cosine term and one has effectively a kind of sine-Gordon theory [52]. It is instructive to discuss this limit where intuition from perturbative QED can be applied.
@@ III Dynamics of expanding strings
Let us now turn to the problem of an expanding QCD string in the reduced model, where the dynamics happens in $d=1+1$ dimensions and massive QED is employed instead of QCD. We start from a quark-antiquark pair with very high energy, where the quark and antiquark move in opposite directions with the speed of light. The electric field that forms between them will be modeled as a coherent scalar field in the bosonized description and we will study its evolution, as well as the excitations around it.
@@ III.1 Coordinate system
Expressed in Bjorken coordinates, the equation of motion for the scalar field as obtained from the variation of Eq. (4) reads [using the ratio in Eq. (7)]
@@ III.2 Background field evolution
In order to fully understand the evolution of the coherent field $\Phi(\tau)$ also at late times, we numerically solve the nonlinear equation of motion, Eq. (10), using the fourth-order Runge-Kutta method. Figure 2 illustrates the behavior of $\Phi(\tau)$ for various choices of the ratio $\kappa$ and the vacuum angles $\theta=0$ as well as $\theta=\pi$.
@@
Note that depending on the value of $\kappa$, there are qualitative differences in the dynamics of the background field. For very small values $\kappa\ll 1$, corresponding essentially to small fermion mass $m\ll q$, the potential has only a single global minimum which is then approached by the classical evolution equation in the asymptotic long time limit. In contrast, when $\kappa$ is larger, the background field can be trapped in a local minimum, which corresponds to a metastable state.
In Sec. V, we will discuss the production of quantum excitations around the coherent background field solution $\Phi(\tau)$. However, before that, we first address the metastable background solutions and the decay of such false vacuum states through field tunneling.
@@ IV False vacuum decay
It is an interesting problem to follow the time evolution in real time within the future light cone of the breaking point at the coordinate origin. This leads to a very similar problem to the one discussed in the previous section. Directly on the light cone, initial conditions are set through the bounce solution, but everything in the future of that is to be computed dynamically.
@@
Displayed on the lower plots of Fig. 3, we show the resulting background field $\Phi$ evolution for the new initial conditions (for $\theta=0$ in the left panel and $\theta=\pi$ in the right panel). Again, the field shows damped oscillations, now around the true vacuum.
@@
Combining both solutions, the analytically continued bounce solution for $|x|>t$ and the solution to classical evolution equations for $t>|x|$ is shown in Fig. 4. The field decays from the false vacuum outside the light cone to the true one inside it, showing damped oscillations very similar to those obtained in Sec. III.2.
@@ V Particle production
Now that we have obtained nontrivial evolving background field configurations in two closely related scenarios, it is interesting to study quantum fluctuations around them. Field excitations can be seen as particles, and in this section, we aim to compute the resulting particle production [63, 64, 65, 66].
@@ V.1 Adding quantum fluctuations
such that the equal-time canonical commutation relation reads with the conjugate momentum field $\Pi(\tau,\eta)=\delta\Gamma/\delta(\partial_{\tau}\phi(\tau,\eta))=\tau\partial_{\tau}\phi(\tau,\eta)$,
@@ V.2 Initial choice of mode functions
Another interesting state to be considered is a coherent state that basically agrees with the vacuum state, but where the expectation value is shifted, $\Phi_{\text{vac}}\to\Phi(\tau_{0})$. We denote this state by $|\Omega,\Phi(\tau_{0})\rangle$ with $\Phi(\tau_{0})$ the field expectation value $\Phi(\tau)$ at the initial time $\tau=\tau_{0}$. For the scenario discussed in Sec. III this initial value is given in Eq. (11), while for the scenario discussed in Sec. IV it is specified in Eq. (17). For the mode functions in this state, we use at very early times $\tau\to 0^{+}$ the form given in Eq. (23) without any modification. Of course, deviations will occur at later times $\tau>0$, see below.
The “shifted vacuum” coherent state specified through this prescription is only one out of many possible quantum states one could consider at $\tau=0^{+}$. In particular for the bounce scenario discussed in Sec. IV one may expect that one could do better and actually determine the quantum state of fluctuations after the bounce possibly through analytic continuation from a calculation in Euclidean space, but we leave this for future work.
Starting with the initial conditions (23) for very early times, we use the differential equations (20) to find solutions for later times. Figure 5 shows the real and imaginary part of the mode functions obtained in this way for different choices of the coupling constant $\kappa$ and the rapidity wave number $k$. In the right panel of Fig. 6, we show the growth in time of the occupation numbers [see Eq. (29) below for their definition] for several wave number $k$ for $\theta=0,\kappa=4.0$.
In general, the linearization procedure and the Bogoliubov theory are valid as long as the effect of the fluctuations effect on the background field is small and if the fluctuations remain small during the evolution. Due to the effect of the Bjorken expansion, the occupation numbers exhibit exponential growth only for a time window, after which their value saturates. This phenomenon is discussed in detail in Ref. [67]. Even though the backreaction of the fluctuations is expected to introduce some extra damping in the background field dynamics, it does not significantly alter the qualitative dynamics governing the mode function equation. This analysis is left for future work.
@@ V.3 Asymptotic particle number
As a consequence of the overall expansion and dilution of energy, the background field $\Phi(\tau)$ eventually approaches a constant asymptotic value for very late Bjorken times, $\Phi(\tau)\to\Phi_{\text{asym}}$ for $\tau\to\infty$. We concentrate on situations where this value agrees with the global minimum, $\Phi_{\text{asym}}=\Phi_{\text{vac}}$. [For $\theta=\pi$, there are two degenerate global minima, and $\Phi(\tau)$ dynamically selects one.]
@@
Using the usual Bogoliubov theory arguments (see, for example, Ref. [48, 68, 69, 70]), one can obtain the following expression for the occupation number of the asymptotic state
@@
Because $f_{k}^{\text{asym}}(\tau)$ is only a proper solution to the mode equation (20) at asymptotically late time, the expression in Eq. (29) only has strict physical meaning in the limit $M\tau\gg 1$. The left panel of Fig. 6 shows the particle number $n_{k}$ against the wave vector $k$ at different times $\tau$ for the choice of parameters $\theta=0$ and $\kappa=4.0$ (left panel). One observes that around $M\tau=510$, the convergence with $M\tau$ is approaching, at least for some choices of $\theta$, $\kappa$, and $k$. The spectrum displays a peak at $k\approx 0$. This is the expected behavior: excitations with small energy can be produced, while for large $k$, the kinetic term $\sim k^{2}$ in Eq. (20) dominates over the time-dependent term in the second line and particle production is suppressed.
@@ V.4 Total particle number and temperature
which are of the same order of magnitude as the hadronization temperature $T_{H}\approx 160$--$170$ MeV [4]. The parameter $\theta=\pi$ seems to be closer to the actual result. However, $\theta=0$ is of the same order of magnitude, indicating that other angles may yield comparable results. Overall, it can be assumed that the numerical outcome is relatively consistent across various angles and different choices of the initialization of the fluctuations and that this result remains robust. Some discrepancies in the results are expected since QED1+1 is only a toy model for QCD. The Schwinger model offers valuable insights into confinement and thermalization mechanisms, suggesting that dimensional reduction and breaking of the confining string through tunneling present in the dynamics of the Schwinger model are key effects for the observation of the apparent thermal features. However, it cannot fully replicate QCD in higher dimensions. It also lacks some important features of QCD, like the gauge dynamics, multiple fermion flavors, and the non-Abelian symmetry group $SU(3)$.
The above results have been obtained for fluctuations around a decaying false vacuum as described in Sec. IV. The calculation can also be done for fluctuations around the background solution described in Sec. III with a similar result, but temperatures given by $T_{\theta=0}\approx 1\text{ GeV},T_{\theta=\pi}\approx 2\text{ GeV}.$
@@ VI.1 Experimental setup
The two components in the system are represented by the field operators $\psi_{1}$ and $\psi_{2}$. Both have the same mass $m$, and their interactions are determined by two coupling constants, namely $h$, the coupling of each component with itself, and $h_{c}$, which is related to an intercomponent interaction. The external potential of the trap in which the atoms are confined has been omitted here. Additionally, an external field, e.g., a radio frequency field, induces the transition between the two states with a transition rate $\nu$. Furthermore, the transition rate $\nu_{0}$ is modulated with radio frequency $\omega$, the amplitude of the modulation being $\nu(t)=\nu_{0}+\delta\omega\cos(\omega t)$. To bring the Hamiltonian (35) into a more convenient form, the fields are expressed in terms of their densities $\rho_{j}=\psi_{j}^{\dagger}\psi_{j}$ and phases $\vartheta_{j}$,
@@
Additionally, we can switch to the set of variables that consists of the mean $\rho$ and the relative density $\epsilon$ defined as $\rho=(\rho_{1}+\rho_{2})/2,\,\epsilon=(\rho_{2}-\rho_{1})/2,$ and the sum $\zeta$ and the difference $\phi$ of the phases $\zeta=\vartheta_{1}+\vartheta_{2},\,\phi=\vartheta_{2}-\vartheta_{1}.$ The relative phase can be measured with an interferometer [76]. We will focus on the relative phase $\phi$ in the following. This should not be confused with the above-mentioned scalar field $\phi$. Moreover, the frequency variation $\omega$ is assumed to be large in comparison to the timescales of the system, specifically $\omega_{0}=2\sqrt{\nu_{0}h\bar{n}}$ where $\bar{n}$ is the expectation value of the mean particle density. This allows time averaging and gives the final effective Lagrangian after integrating out the densities $\rho$ and $\epsilon$